OpenStudy (anonymous):

Use Stokes' Theorem to calculate the circulation of the field F around the curve C when F and C are the following: F= (x^2)i+(2x)j+(z^2)k C= 4x^2+y^2=4 in the xy-plane

OpenStudy (anonymous):

find curl of F= (x^2)i+(2x)j+(z^2)k

OpenStudy (anonymous):

And how do I do that?

OpenStudy (anonymous):

So i take the cross product of the individual components of the vector and the partial derivatives to find the curl. That I get. Where do I go from there?

OpenStudy (anonymous):

get a surface , any surface whose boundary curve is C

OpenStudy (anonymous):

So C is given by the problem. Then what?

OpenStudy (anonymous):

you have to find a surface (an easy one) whose boundary curve is C

OpenStudy (anonymous):

4x^2+y^2=4 is a ellipse

OpenStudy (anonymous):

follow this http://tutorial.math.lamar.edu/Classes/CalcIII/StokesTheorem.aspx

OpenStudy (anonymous):

So I need to parametrize C first...but I'm not really sure how to do that.

OpenStudy (anonymous):

is it [r=4\cos \theta I + \sin \theta J + K\]

OpenStudy (anonymous):

r=4cos(0)i + sin(0)j + k where 0 = theta

OpenStudy (anonymous):

hold on, let me think it through

OpenStudy (anonymous):

r=cos(t)i+2sin(t)j+k F(r(t))=cos^2(t)i+2cos(t)j+k dr/dt=-sin(t)i+2cos(t)j F(r(t)).(dr/dt)=-sin(t)cos^2(t)+4cos^2(t) integration from 0 to 2pi of this gives 4pi